How many ways are there of choosing 5 letters, without replacement, from 17 distinct letters, if
the order of the choices is not taken into consideration?
the order of the choices is taken into consideration?

Answer by stanbon(57275)

How many ways are there of choosing 5 letters, without replacement, from 17 distinct letters, if
the order of the choices is not taken into consideration?
Ans: 17C5 = [17!/(17-5)!]/5! = 6188
---------------------------
the order of the choices is taken into consideration?
17P5 = 17!/(17-5)! = 17*16*15*14*13 = 742,560
----------------------------------------------
Cheers,
Stan H.